Institut for Matematik og Datalogi Odense Universitet January 30, 1996

JFB

DM11 – F96 – Lecture 1

Textbook

Kenneth H. Rosen, Discrete Mathematics and Its Applications, Third edition, McGraw-Hill, Inc., 1995. Note that answers to the odd-numbered exercises can be found at the back of the book. In addition, some concepts used in the exercises are defined between exercises, so when a problem uses a concept or notation not defined in the text, you should look to see if it is defined a bit before the problem in question. The following topics which will be covered in the course are not adequately covered in the text, so there will be brief notes about them:

• structural induction

• some algebra, especially basic group theory

• Huffman coding

Format

There will be weekly assignments (which are not obligatory) which you should be prepared to discuss in the discussion sections (examinatorier) the week after the problems are assigned. The instructors are Per Vikkelsøe and Anders Yeo. The sections will be held in U10 and U13 at 12:15 on Thursdays. (Note that neither section will meet in week 6. Both begin in week 7.) Some of the problems on some of the assignments will be marked indicating that if you give your instructor your solution to the problem, he will look at and give you comments a week later. Thus, you have the opportunity of getting advice on your proof-writing style. There will be a written exam in June, with questions written in Danish (you can write it in Danish or English). The weekly notes and other information about the course is available through the WorldWideWeb. Use the URL:

http://www.imada.ou.dk/ joan/dm11/index.html

Lecture, February 7

We begin with an introduction to the course. Then, we will cover sections 1.1 through 1.5 in the textbook.

Lecture, February 14

We will finish chapter 1, but we will skip section 1.8 which is covered in DM02.

Problems to be discussed in week 7

7.a, 8.b, 8.c, 10.b, 10.c, 26, 27, and 28 from section 1.2. 21, 26 and 30 from section 1.3. 12 and 14 from section 1.4. Problem 10.b from section 1.2 is probably the most challenging.